### v1

# Solving a Stackelberg Game on Transportation Networks in a Dynamic Crime Scenario: A Mixed Approach on Multi-Layer Networks

Abstract
Interdicting a criminal with limited police resources is a challenging task as the criminal changes location over time. The size of the large transportation network further adds to the difficulty of this scenario. To tackle this issue, we consider the concept of a layered graph. At each time stamp, we create a copy of the entire transportation network to track the possible movements of both players, the attacker and the defenders. We consider a Stackelberg game in a dynamic crime scenario where the attacker changes location over time while the defenders attempt to interdict the attacker on his escape route. Given a set of defender strategies, the optimal attacker strategy is determined by applying Dijkstra’s algorithm on the layered networks. Here, the attacker aims to minimize while the defenders aim to maximize the probability of interdiction. We develop an approximation algorithm on the layered networks to find near-optimal strategy for defenders. The efficacy of the developed approach is compared with the adopted MILP approach. We compare the results in terms of computational time and solution quality. The quality of the results demonstrates the need for the developed approach, as it effectively solves the complex problem within a short amount of time.
keywords:
Resource allocation, Multi-layer time expanded network, Vehicle routing and scheduling, Network optimization, Exact algorithm, Approximation algorithm
1 Introduction
We consider a Stackelberg game where the defender is the leader, and the attacker is the follower. The assumption is that the defenders have information only about the crime location in a large transportation network. In this Stackelberg game, first, the defenders commit to a strategy, and then the attacker generates the best response to the given defender strategy. We use the concept of a multi-layer network in which, at each time-stamp, we create a copy of the entire network. We consider edge length as the time factor to create connections between these multi-layer networks. This means that depending on the edge length, we choose the layers from which the start and end nodes of that particular edge are selected. In this way, for all edges in the original transportation network, we create corresponding edges in the multi-layer network. We consider a set of strategies for both players, i.e., the defenders and the attacker. For the defender, we use a mixed strategy set where each strategy is assigned a mixed probability, with the sum of these probabilities equaling one. For the attacker, we consider a pure strategy set.
To generate the optimal attacker strategy, we first assign weights to the nodes depending on the mixed probabilities and the nodes’ presence in the defender strategies. Considering all defender strategies, we sum up the probabilities if any node is present in multiple defender strategies. Then, we assign the weight of each node to all of its incoming edges. Finally, we apply Dijkstra’s algorithm on this network to find the optimal strategy for the attacker. For the defender, we create a multi-layer network in the same way as for the attacker. Weights are assigned to all nodes based on the attacker strategies, with each attacker strategy given equal probability. We develop a novel approximation algorithm for the defender to generate a near-optimal defender strategy. Thus, for the defender, the problem can be considered as finding a near-optimal path with the aim of covering at least one node from each attacker strategy. For this, we use different colors to represent each attacker strategy. The developed defender strategy includes as many different colored vertices as possible. We consider a Stackelberg game instead of a zero-sum game. The advantage of the Stackelberg game over the zero-sum game is that it allows for a more realistic representation of real-world scenarios where one player has more information, resources, or power than the other player. In this Stackelberg game, the defenders have an advantage because they can consider the attacker’s potential moves and responses when making their decisions. This can result in a more efficient outcome than in a zero-sum game, where the players are equally matched.
Due to the complex transportation network and limited police resources, generating efficient strategies for both players is a challenging task. We can summarize the main contributions of this paper as follows: (1) To the best of our knowledge, this is the first attempt to solve a Stackelberg game considering the escape interdiction game (EIG) problem using the multi-layer network concept, with a focus on developing solution methodologies. (2) We develop an optimal approach for the attacker and a near-optimal approach for the defenders, and we are not aware of any previous implementations for escape interdiction problems on multi-layer network. (3) We generate a multi-layer network to create efficient strategies for both players. (4) We use grids dataset for the case study and show that our MLN-EIGS approach outperforms MILP-EIGS approach in terms of solution quality, with a minimal performance gap.
The paper is organized as follows. In Section 2, we present the relevant research. In Section 3, we define the problem description and modeling. In Section 4, we describe the solution methodology. Section 5 presents the quality of the results. We conclude this research in Section 6.
2 Related Work
Security games play an important role in providing social security. Zhang et al. [3] consider a zero-sum game for the escape interdiction problem and develop a MILP-based solution approach to provide an optimal solution. To demonstrate the efficacy of their methodology, they generate optimal solutions on grids of different sizes. Samanta et al. [1] consider the same zero-sum game problem and develop a meta-heuristic-based solution approach to provide a scalable near-optimal solution in a time-efficient manner. For the same escape interdiction problem, Samanta et al. [2] develop a simulation-based approach to generate a scalable solution to increase security in a large transportation network.
Recent literature focuses on Stackelberg games considering the escape interdiction problem to decrease the crime rate in society. For example, Zychowski and Mandziuk [4] consider search games (SEG) on directed graphs. They consider multiple defender resources and one attacker, where the attacker aims to reach one of several predefined target points from a fixed location. They develop a genetic algorithm-based heuristic approach to provide a near-optimal solution on synthetic datasets.
Basilico et al. [9] consider a leader-follower game and formulate the problem as a mathematical programming model. They use optimization software tools to solve the model by generating the optimal strategies for both leader and follower. Letchford and Conitzer [11] consider security games on graphs and develop a polynomial-time algorithm to generate optimal strategies for players. Similarly, Iwashita et al. [10] consider security games on graphs and develop an algorithm to reduce the graph by eliminating unnecessary edges, providing a time-efficient, scalable near-optimal solution. In the same vein, Karwowski and Mandziuk [18] consider an evader-defender Stackelberg game model and develop a Monte Carlo Tree Search approach to provide efficient patrolling schemes.
Tsai et al. [13] consider an attacker-defender Stackelberg game model and develop a linear program to generate optimal mixed strategies for players by allocating limited resources optimally, with a case study on part of the Mumbai road network. Similarly, Shieh et al. [12] develop a game-theoretic system to provide security with limited resources in the port of Boston and also test its efficacy in the port of New York. They schedule patrols efficiently to provide optimal mixed strategies for players, considering an attacker-defender Stackelberg game model. Similar papers focus on Stackelberg security games to provide security in society (e.g., Cermak et al. [16], Lou et al. [19], Sinha et al. [14], ˇCern‘y et al. [17], Zhang and Malacaria [15]).
Conitzer and Sandholm [5] show that exact approaches are not suitable because of the NP-hard nature of this problem, as it faces scalability issues due to the high time complexity. Paruchuri et al. [6] develop a MILP-based exact algorithm to solve Bayesian Stackelberg security games. Bosansky and Cermak [7] also consider Stackelberg games and develop algorithms based on LPs and MILP to generate Strong Stackelberg Equilibrium (SSE) and perform a case study at Los Angeles International Airport, considering security scheduling.
The layered graph concept is a useful tool for solving complex problems on transportation networks in a time-efficient manner. For example, Saito et al.[8] consider the problem of selecting important nodes in a network and construct a layered graph from the original graph, where each layer is added on top as time proceeds to demystify the complex problem. Hence, we focus on solving a Stackelberg escape interdiction game using the layered graph concept.
3 Problem Description
We consider a two-player Stackelberg game where this sequential game is played between multiple defenders $(D)$ and one attacker $(A)$. The defenders have a finite set of actions $\acute{S}$, and the attacker has a finite set of actions represented by $\acute{A}$. $S$ and $A$ represent the sets of probability distributions over the defender and attacker actions, respectively. The game is modeled as a general-sum game, where $x\in S$ and $y\in A$ are the probability distributions over $\acute{S}$ and $\acute{A}$, respectively.
For $a_{j}=(v_{j},t^{a}_{j})$ and $s^{r}_{i}=(v^{r}_{i},t^{r,in}_{i},t^{r,out}_{i})$, $d_{r}$ reaches the attacker at node $v^{r}_{i}$ if $v^{r}_{i}$ = $v_{j}$ and $t^{r,in}_{i}\leq t^{a}_{j}\leq t^{r,out}_{i}$. The defender gets -1 if the attacker successfully escapes from the transportation network; otherwise, the defender’s utility value is 0 (see Eq. 1). Considering this case, we define $Z_{s^{r}_{i},a_{j}}=1$, otherwise $Z_{s^{r}_{i},a_{j}}=0$.
Here, $y$ and $A$ are used to calculate the expected utility of the defender, which is represented in Eqs. (2) and (3).
$U_{d}\;(S,A)=\begin{cases}0,&\text{if $\exists$$Z_{s^{r}_{i},a_{j}}$ = 1, $a_{j}$ $\in$ A, $s^{r}_{i}$ $\in$ $S^{r}$, r $\in$ R}.\\ -1,&\text{ otherwise}.\end{cases}$ (1)
$U_{d}\;(x,A)=\sum\limits_{S\in\textbf{S}}U_{d}(S,A)\;x_{s}$ (2)
$U_{d}\;(x,y)=\sum\limits_{A\in\textbf{A}}U_{d}(x,A)\;y_{A}$ (3)
Therefore, we need an optimization model so that the defenders can find their best optimal strategy by solving the following linear problem (see Eqs. 4-6).
$max\qquad U^{*}$ (4)
$s.t.\qquad U^{*}\leq U_{d}\;(x,A)\qquad\forall A\in A^{{}^{\prime}}$ (5)
$\sum\limits_{S\in\textbf{S}}x_{s}=1,x_{s}\geq 0\qquad\forall S\in S^{{}^{\prime}}$ (6)
The expected utility of the defender and attacker are $U_{d}$ and $U_{a}$, respectively. For any pair of strategies $(x,y)$, the expected payoff of the defender is denoted by $U_{d}(x,y)$.
First, in Eq. (7), the defender commits to a strategy $x\in X$, and given such an $x$, the attacker chooses his strategy from the best-response set $BR(x)$ where
$BR(x)=\;argmax_{y\in Y}\;U_{a}\;(x,y)$ (7)
Second, in Eq. (8), to maximize utility, the defender chooses the strategy $x$ (the best response of the defender) to the attacker’s best response.
$max_{x\in S}\;U_{d}\;(x,y)\qquad s.t.\;y\in BR(x)$ (8)
Third, in a Strong Stackelberg Equilibrium (SSE), we assume that the follower (attacker) breaks ties in favor of the leader (defender) (see Eq. 9). In that case, the optimization problem is
$max_{x\in S,\;y\in BR(x)}\;U_{d}\;(x,y)$ (9)
SSE is popular because it is always guaranteed to exist ((Conitzer [20]).
4 Proposed Mixed Approach on Multi-Layer Networks
This section presents the proposed methodology for the considered escape interdiction problem which is formulated as a Stackelberg game model. A novel mixed approach on layered graph concept is developed to solve the Stackelberg model, where the defender is the leader and the attacker is the follower. For the defender, the developed approximation algorithm finds the efficient movement of the defender from a random initial location. For the attacker, the initial strategy is any random path from the crime location to any exit point. An exact algorithm is developed to generate the optimal path for the attacker using Dijkstra’s algorithm on the time-expanded network. The new defender strategy is generated using a greedy algorithm on the time-expanded network. The defender’s mixed strategy set is updated based on the recalculated utility. The defender then selects the best strategy to commit to. This best defender strategy is given as input to the attacker. The attacker tries to generate a new optimal strategy using the exact approach, taking the new defender strategy as input. The attacker updates the pure attacker strategy set, where all strategies are of equal probability. The algorithm checks if, for $K$ consecutive iterations, the game utility is below a threshold value named $\epsilon$. The value of $\epsilon$ is set to 0.05 for our experiments. If this convergence criterion is met, the algorithm converges, and we return the best defender strategy with the probability of interdiction and total runtime (see Algorithm 1). The flowchart is shown in Fig. 1.
Figure 1: Flowchart of Stackelberg Game Model
Input: Initialize the initial strategy sets of defender and attacker $S^{\prime},A^{\prime}$;
Output: $U_{d}:$ Defenders’ game utility;
repeat
       $(x,y)\leftarrow CoreLP(S^{\prime},A^{\prime})$;
       Defender commits to the best strategy $x^{\star}\in x$ which provides maximum utility of defender ;
       $BR:\;A^{*}\leftarrow{ExactAO(x^{\star})}$;
       if $A^{*}\neq\emptyset$ then
             $A{{}^{\prime}}\leftarrow A^{{}^{\prime}}\cup A^{*}$;
       end if
      $S^{*}\leftarrow{ApproxDO(A^{{}^{\prime}})}$;
       if $S^{*}\neq\emptyset$ then
             $S^{{}^{\prime}}\leftarrow S^{{}^{\prime}}\cup S^{*}$;
       end if
until convergence;
return $U_{d}$.
Algorithm 1 MLN-EIGS Algorithm
To create a benchmark, we construct a novel Stackelberg game where the defender is the leader, and the attacker is the follower. The best strategies for the attacker and defender are generated using bestAO and bestDo, which are the MILP approaches for the attacker and defender, respectively (see Algorithm 2). Zhang et al. [3] develop these optimal approaches to provide optimal strategies for attackers and defenders given a strategy set for each player. Similar to the MLN-EIGS, in this MILP-EIGS the defender commits to the best mixed strategy. The attacker takes this best defender mixed strategy as input, generates the optimal attacker strategy, and updates the pure attacker strategy set. Then, we calculate the game utility, and the game continues until the game utility is below $\epsilon$ for $K$ consecutive iterations. Here $\epsilon$ acts as a threshold with a value of 0.05. We return the defender utility value once the convergence is reached.
Input: Initialize the initial strategy sets of defender and attacker $S^{\prime},A^{\prime}$;
Output: $U_{d}:$ Defenders’ game utility;
repeat
       $(x,y)\leftarrow CoreLP(S^{\prime},A^{\prime})$;
       Defender commits to the best strategy $x^{\star}\in x$ which provides maximum utility of defender ;
       $BR:\;A^{*}\leftarrow{BestAO(x^{\star})}$;
       if $A^{*}\neq\emptyset$ then
             $A{{}^{\prime}}\leftarrow A^{{}^{\prime}}\cup A^{*}$;
       end if
      $S^{*}\leftarrow{BestDO(A^{{}^{\prime}})}$;
       if $S^{*}\neq\emptyset$ then
             $S^{{}^{\prime}}\leftarrow S^{{}^{\prime}}\cup S^{*}$;
       end if
until convergence;
return $U_{d}$.
Algorithm 2 MILP-EIGS Benchmarking Algorithm
4.1 Efficient Attacker Strategy Design Using Exact Approach on Time Expanded Network
The optimal attacker strategy is generated using the exact approach for the attacker (see Algorithm 3). For each time-stamp, we create a copy of the entire network. Then, we establish connections in this multi-layer network using the edge lengths in the original graph. This algorithm describes how we assign the probability of interdiction to each node. Finally, we apply Dijkstra’s algorithm on this multi-layer network to generate the optimal strategy for the attacker.
Input: Crime node is the “START” node, all exit nodes are the “GOAL” nodes, Original graph $(G_{0})$;
Output: Best attacker strategy having minimum probability of interdiction;
Construction of layered graphs:
for $(i=0;i<t_{max};i++)$ do
       Generate one copy of the original graph/network, labelled as $G_{i+1}$ ;
end for
Connect the Layered Graphs depending on time/distance between adjacent nodes in the original graph:
for all edges in the original graph do
       $L$ = current edge length, $S$ = From node, $T$ = To node (in original graph) ;
       for $(j=0;j<t_{max}-L;j++)$ do
             Select graph $G_{j}$ and graph $G_{j+L}$ ;
             Create edge from node $S$ of $G_{j}$ to node $T$ of $G_{j+L}$ ;
       end for
end for
Update the minimum Prob. of. Interdiction $(1-P)$ value to the node depending on its presence in the defender strategy:
Initial probability of interdiction $(P)$ of all nodes in all layered graphs = 0 ;
for all $N$ defender strategies in the best strategy set do
       for all nodes present in the current strategy do
             $Node=n,t_{in}=In\;Time,t_{out}=Out\;Time$ while $t_{in}<t_{out}$ do
                   Select node $n$ of graph $G_{t_{i}n}$ Update the minimum Prob. of Interdiction of that node $n:(P=P+P_{mix})$ where $P_{mix}$ is the mixed prob. of the current defender strategy. $t_{in}=t_{in}+1$
             end while
       end for
end for
Algorithm 3 Optimal Attacker Strategy Design Using Time Expanded Network
Update the minimum Prob. of. Interdiction $(1-P)$ value to the node depending on its presence in the defender strategy:
for all nodes present in the layered graph do
       if Prob. of interdiction of a node $n$ is $P$ then
             Assign the weight $P$ to all in-coming edges of that node.
       end if
end for
Apply Dijkstra’s algorithm on the developed Multi-Layer Network.
return $U_{d}$.
4.2 Efficient Defender Strategy Design Using Approximation Algorithm on Time Expanded Network
We develop a novel approximation algorithm to generate a near-optimal strategy for the defender in a time-efficient manner. We describe the approach in the following steps. Since the attacker’s strategy set consists of pure strategies, we assign different colors to each of these attacker strategies. Then, we attempt to construct a path for the defender that includes at least one colored vertex to interdict each attacker strategy. This implies that the defender’s strategy contains as many different colored vertices as possible.
  • 1.
    Input
    • (a)
      A directed acyclic weighted graph, where some vertices are colored.
    • (b)
      There are $k$ different colors $(c_{1},c_{2},...,c_{k})$.
    • (c)
      A threshold value of the path length $t$.
    Here, each colored vertex corresponds to a particular attacker’s strategy (a path).
  • 2.
    Goal:
    To find a path whose length is at most $t$ and contains as many different colored vertices as possible.
  • 3.
    Guess:
    This problem is difficult as Dijkstra cannot solve its decision version, which is $NP$-complete. We provide a formal proof in this paper.
We can use the steps below for the approximate algorithm.
  • 1.
    Step1: Find a shortest path to any one of colored vertices from the start vertex (we can use Dijkstra). Assume the path is to vertex $v_{1}$ with color $c_{i_{1}}$.
  • 2.
    Step2: Find a shortest path to any one colored vertices except $c_{i_{1}}$ from $v_{1}$. Assume the path is to vertex $v_{2}$ with color $c_{i_{2}}$.
  • 3.
    Step3: Find a shortest path to any one colored vertices except $c_{i_{1}}$ and $c_{i_{2}}$ from $v_{2}$, and so on, until all colors are visited or the total path length reaches $t$.
If we cannot find a path that covers all colors, we can create a tree (or a set of paths) instead, which corresponds to a non-deterministic strategy for a defender.
In a simplest form, if the path we obtained in the previous method does not cover a subset of colors, we create another path from the initial vertex, which tries to cover these remaining colors only. We repeat this procedure until all colors are covered. Then, the defender flips a coin and chooses one path.
4.3 Theorem: Problem ”color covering” is $NP$-complete
Problem ”color covering”
  • 1.
    Input: A directed graph and the initial vertex. Some vertices are colored. There are $m$ different colors. It is possible that one vertex has multiple colors.
  • 2.
    Output: ”yes” when there exists a path from the initial vertex with length $n$, such that all $m$ colors appear at least one of a vertex on the path. ”no” otherwise.
Theorem: Problem ”color covering” is NP-complete
  • 1.
    Proof idea: Reduction from 3-SAT (which is known to be NP-complete).
Problem 3-SAT
  • 1.
    Input: $n$ boolean variables $(x_{1},\ldots,x_{n})$, $m$ clauses. Each clause is a disjunction of three literals. Each literal is a variable or its negation.
  • 2.
    Output: ”yes” if there exists an assignment of variables that makes all clauses true. ”no” otherwise.
For a given $3-SAT$ instance, we create an $n+1$ level network.
  • 1.
    There is one level-00 vertex, which is the initial vertex.
  • 2.
    There are two level-$i$ vertices (for $i>0$). One vertex corresponds to making variable $x_{i}$ true. The other vertex corresponds to making variable $x_{i}$ false.
  • 3.
    There exists a directed edge from each of level-$i$ vertex to each of level-$(i+1)$ vertices.
  • 4.
    Each clause has its own color.
  • 5.
    If a clause with color $c$ contains $x_{i}$, the ”true” vertex for $x_{i}$ has color $c$.
  • 6.
    If a clause with color $c$ does not contain $x_{i}$, the ”false” vertex for $x_{i}$ has color $c$.
Thus, the $3-SAT$ instance is satisfiable iff there exists a path from the initial vertex with length $n$, which covers all colors.
5 Results and Discussion
In this section, we present the results of the developed approaches. The proposed algorithms are coded in Python 3.6 and tested on a computer equipped with an Intel(R) Core(TM) 3.20 GHz processor and 8 GB RAM, operating under the LINUX environment. All MILPs are solved using CPLEX (version 12.8).
In Fig. 2, we consider a sample network and use Dijkstra’s algorithm on the time-expanded network to generate the optimal attacker strategy. Here, we provide two mixed defender strategies as input with probabilities of 0.6 and 0.4. We demonstrate that our developed exact approach for the attacker can generate the optimal attacker strategy, enabling the attacker to escape without interdiction in a concise amount of time (see Table 1).
Figure 2: Optimal Attacker Strategy Design Using Dijkstra Algorithm on Time Expanded Network.
Table 1:
Optimal Attacker Strategy Design Using Dijkstra Algorithm on Time Expanded Network.
Game Parameters
Network Size: Crime Node: $0\_1$, Police Station: $0\_6$, $T_{max}$: 6, Exit Point: 5
Test Case Optimal Attacker strategy Utility of the Final Optimal Attacker Strategy Run Time (Sec)
1 [$0\_1,3\_3,5\_4$] 0.0 0.004
2 [$0\_1,3\_3,5\_4$] 0.0 0.004
3 [$0\_1,3\_3,5\_4$] 0.0 0.0039
4 [$0\_1,3\_3,5\_4$] 0.0 0.004
5 [$0\_1,3\_3,5\_4$] 0.0 0.004
6 [$0\_1,3\_3,5\_4$] 0.0 0.004
7 [$0\_1,3\_3,5\_4$] 0.0 0.004
8 [$0\_1,3\_3,5\_4$] 0.0 0.004
9 [$0\_1,3\_3,5\_4$] 0.0 0.004
10 [$0\_1,3\_3,5\_4$] 0.0 0.004
In Fig. 3, we consider a sample network and use the approximation algorithm on the time-expanded network to generate the near-optimal defender strategy. Here, we provide three pure attacker strategies as input with equal probability. We demonstrate that our developed approach for the defender can generate an efficient defender strategy that interdicts all attacker strategies quickly (see Table 2).
Figure 3: Defender Strategy Design Using Approximation algorithm.
Table 2:
Defender Strategy Design Using Approximation algorithm.
Game Parameters
Network Size: Crime Node: $0\_1$, Police Station: $0\_6$, $T_{max}$: 6, Exit Point: 5
Test Case Final Defender strategy No. of Non-interdicted Attacker Strategies by the Final Defender Strategy Run Time (Sec)
1 [$0\_6,2\_3,4\_4,5\_5,6\_5$] 0.0 0.0086
2 [$0\_6,2\_3,4\_5,5\_5,6\_5$] 0.0 0.0090
3 [$0\_6,2\_3,4\_5,5\_5,6\_5$] 0.0 0.0081
4 [$0\_6,2\_3,4\_5,5\_5,6\_5$] 0.0 0.0081
5 [$0\_6,2\_3,4\_5,5\_5,6\_5$] 0.0 0.0082
6 [$0\_6,2\_3,4\_4,5\_5,6\_5$] 0.0 0.0080
7 [$0\_6,2\_3,4\_4,5\_5,6\_5$] 0.0 0.0082
8 [$0\_6,2\_3,4\_4,5\_5,6\_5$] 0.0 0.0081
9 [$0\_6,2\_3,4\_4,5\_5,6\_5$] 0.0 0.0083
10 [$0\_6,2\_3,4\_5,5\_5,6\_5$] 0.0 0.0081
We compare the final defenders’ utility generated by the developed MLN-EIGS algorithm with the final defenders’ utility using MILP-EIGS algorithm, which uses the exact approaches named bestDo and bestAo, developed by Zhang et al. [3]. To establish a benchmark, we solve the Stackelberg Game using the bestDo and bestAo exact approaches. We compare both the solution quality and computational time.
We have generated 10 test cases for each grid from 3*3 to 9*9 (see Tables 3, 4, 5, 6, 7, 8, and 9). All test case instances are generated by varying the crime node location, initial location of defenders, $t_{max}$ (maximum run time), and exit node. For each test case, we check the performance gap and the variation in computational time.
Table 3:
Grid: 3*3.
Test Case Final Defender Utility Run Time (Sec)
MILP-EIGS MLN-EIGS MILP-EIGS MLN-EIGS
1 0.0 0.0 3.71 3.59
2 0.0 0.0 3.69 3.49
3 0.0 0.0 3.72 3.38
4 0.0 0.0 3.78 3.52
5 0.0 0.0 3.72 2.16
6 -1.0 -1.0 3.69 2.54
7 0.0 0.0 3.86 2.23
8 0.0 0.0 3.70 2.11
9 0.0 0.0 3.74 2.13
10 -1.0 -1.0 3.73 2.79
Table 4:
Grid: 4*4.
Test Case Final Defender Utility Run Time (Sec)
MILP-EIGS MLN-EIGS MILP-EIGS MLN-EIGS
1 -1.0 -1.0 12.54 0.72
2 -1.0 -1.0 9.28 0.71
3 -1.0 -1.0 12.53 0.70
4 -1.0 -1.0 12.19 0.70
5 -1.0 -1.0 12.22 0.52
6 -1.0 -1.0 13.32 0.54
7 -1.0 -1.0 12.93 0.52
8 -1.0 -1.0 11.98 0.51
9 0.0 0.0 12.23 0.55
10 -1.0 -1.0 13.38 0.92
Table 5:
Grid: 5*5.
Test Case Final Defender Utility Run Time (Sec)
MILP-EIGS MLN-EIGS MILP-EIGS MLN-EIGS
1 0.0 0.0 20.35 7.39
2 0.0 0.0 20.24 9.52
3 0.0 0.0 20.86 7.58
4 0.0 0.0 43.62 2.09
5 0.0 0.0 34.86 2.12
6 -1.0 -1.0 37.86 2.54
7 -1.0 -1.0 34.93 2.48
8 -1.0 -1.0 40.28 1.12
9 -1.0 -1.0 37.98 1.11
10 0.0 0.0 35.12 1.00
Table 6:
Grid: 6*6.
Test Case Final Defender Utility Run Time (Sec)
MILP-EIGS MLN-EIGS MILP-EIGS MLN-EIGS
1 0.0 0.0 72.75 4.14
2 0.0 -1.0 74.52 4.11
3 -1.0 -1.0 72.39 4.02
4 -1.0 -1.0 72.07 3.97
5 -1.0 -1.0 73.27 2.40
6 0.0 0.0 60.17 5.93
7 0.0 0.0 35.11 1.88
8 0.0 0.0 34.42 4.00
9 -1.0 -1.0 60.49 4.62
10 -1.0 -1.0 34.54 4.61
Table 7:
Grid: 7*7.
Test Case Final Defender Utility Run Time (Sec)
MILP-EIGS MLN-EIGS MILP-EIGS MLN-EIGS
1 0.0 0.0 231.86 15.51
2 0.0 0.0 234.26 16.09
3 0.0 0.0 233.18 20.97
4 -1.0 -1.0 140.45 31.64
5 0.0 0.0 80.82 14.76
6 0.0 0.0 138.56 15.29
7 0.0 0.0 205.43 15.54
8 0.0 0.0 81.21 15.14
9 -1.0 -1.0 140.92 15.09
10 0.0 0.0 143.20 14.88
Table 8:
Grid: 8*8.
Test Case Final Defender Utility Run Time (Sec)
MILP-EIGS MLN-EIGS MILP-EIGS MLN-EIGS
1 -1.0 -1.0 437.82 43.97
2 -1.0 -1.0 436.68 38.61
3 0.0 0.0 422.33 30.94
4 0.0 0.0 399.56 37.32
5 -1.0 -1.0 406.86 31.32
6 -1.0 -1.0 305.71 29.39
7 0.0 0.0 260.80 28.47
8 0.0 0.0 440.14 28.58
9 0.0 0.0 288.26 27.74
10 0.0 0.0 420.42 29.52
Table 9:
Grid: 9*9.
Test Case Final Defender Utility Run Time (Sec)
MILP-EIGS MLN-EIGS MILP-EIGS MLN-EIGS
1 0.0 0.0 526.49 48.69
2 -1.0 -1.0 349.04 43.79
3 0.0 0.0 354.68 35.70
4 0.0 -0.5 357.99 42.82
5 0.0 0.0 498.56 44.34
6 0.0 0.0 1069.11 43.36
7 0.0 0.0 735.04 44.10
8 0.0 0.0 795.54 36.17
9 0.0 0.0 846.62 35.70
10 0.0 0.0 577.01 81.17
We show that for these small-size grids, our developed mixed approach named MLN-EIGS can provide an optimal solution like the MILP approach named MILP-EIGS in most cases and takes less computational time compared to MILP-EIGS.
6 Conclusion and Future Directions
In this paper, we consider a Stackelberg game theoretic model to solve security games considering the dynamic crime scenario. This paper proposes a novel optimal routing and resource allocation approach on a transportation network. The optimal strategy for the attacker is generated using an exact approach. To the best of our knowledge, this is the first time an exact approach using the multi-layer network concept is used for the attacker to solve a security game. This developed exact approach consists of Dijkstra’s algorithm and the layered graph concept. The near-optimal strategy for the defender is generated using a novel approximation algorithm and the layered graph concept on a transportation network in a time-efficient manner. The experiments are conducted on 3x3 to 9x9 grids.
The results are compared with the solutions developed using MILP approaches for both attacker and defender in this Stackelberg game scenario. For benchmarking, we formulate the Stackelberg game to provide optimal solutions using the adopted bestAO and bestDO approaches. The contribution of the paper is that the developed MLN-EIGS can offer quality solutions like MILP-EIGS but in a significantly shorter amount of time.
The developed approaches have a few limitations. For example, the proposed approach does not include real-time traffic. Future research can focus on developing efficient algorithms to incorporate traffic into the developed mixed approach. In addition, an exact approach for the defender can also be developed to generate an optimal strategy for the defender on a time-layered network.